. Let the sum of n, 2n, 3n terms of an A.P. be S1,S2and S3 respectively, show that : S3 = 3(S2 - S1).

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student
we know that\sum n=\frac{n(n+1)}{2}….......(1)
similarly 2\sum n=2\frac{n(n+1)}{2}….........(2)
and3\sum n=3\frac{n(n+1)}{2}…...............(3)
From 1,2 and 3 it is clear that s3=3(s2-s1)
Thanks and regards
Shaik Aasif
askIITians faculty